### v1

@@ 4.3 Complexity Analysis
For a fixed dimension $d$, we assume that there are $n$ hyperplanes, $v$ vertices in the arrangement, and $v_{d}$ vertices whose associated dictionary needs to apply the Zero rule $d$ times to arrive. When calculating a new dictionary, one needs $(n-d)(d+1)=\mathcal{O}(nd)$ calculations. Testing the lex-min also needs $\mathcal{O}(nd)$ calculations.
The computational complexity of the VE algorithm using the Criss-Cross rule ideally is only $\mathcal{O}(n^{2}dv)$, which has been provided in [4]. However, the rule used in [4] does not fulfill the if-and-only-if condition, which means that it may trap into loops when encountering degenerate vertices or skip certain vertices. We have constructed an example to illustrate this point later. Thus, the actual computational complexity is far worse than $\mathcal{O}(n^{2}dv)$ in some cases.
@@
Now, let us analyze the complexity of the method from Moss [25]. In each dictionary, one needs to compute the ratio of the column of $x_{g}$ with every column and then select the extreme value. This step incurs a cost of $\mathcal{O}(d(n-d)v)=\mathcal{O}(ndv)$. Subsequently, during the comparison with discovered dictionaries, their co-basis need to be checked. If there are $k$ recorded dictionaries, then comparison requires $\mathcal{O}(kd)$ for each candidate. During the procedure of generating the adjacent co-basis about a fixed dictionary, one may find at most $nd$ candidates from it, comparing each candidate takes $\mathcal{O}(knd^{2})$ computations. Since there are a total of $\mathcal{O}(v)$ dictionaries, the total amount of computations is $\mathcal{O}(\sum_{k=1}^{v}(knd^{2}))=\mathcal{O}(nd^{2}v^{2})$. It’s important to note that one only needs to compute those dictionaries whose co-basis are not duplicated, which costs $\mathcal{O}(ndv)$ in computing dictionaries and determining if they pass the lex-min test. Consequently, the complexity of the method from Moss [25] in VE is $\mathcal{O}(ndv+nd^{2}v^{2}+ndv)=\mathcal{O}(nd^{2}v^{2})$.
@@
As for the complexity of our algorithm, at each entry of a dictionary, testing a valid reverse of the Zero rule requires at most $\frac{1}{2}(n-d+n)d$ examination, which is in $\mathcal{O}(nd)$. $\mathcal{O}(v-v_{d})$ dictionaries need to test the validity, while $\mathcal{O}(v_{d})$ do not need. Therefore, the complexity is approximately given by
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d}).$ (8)
@@
The proportion of vertex $v_{d}$ in the overall vertex has a significant impact on the actual complexity. Below, we estimate $v_{d}$ by considering two cases: one is a special case, and the other is a simple arrangement. Before estimation, we first characterize what kinds of dictionaries lie in the $d$-th layer.
@@
In the case of parallel hyperplanes, where hyperplanes form boundaries of hypercubes arranged along a specific coordinate axis, such as $y_{1},y_{2},\ldots,y_{d}=0,1,y_{1}=2,y_{1}=3,y_{1}=4,\ldots,y_{1}=n-2d+1$. As shown in the LHS of Figure 8, $v_{d}=\frac{v}{2^{d}}$, then Eq. (8) can be simplified to
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})=\mathcal{O}((1-\frac{1}{2^{d}})n^{2}d^{2}v+\frac{1}{2^{d}}ndv)=\mathcal{O}((1-\frac{1}{2^{d}})n^{2}d^{2}v).$
If one stacks multiple layers of hypercubes on top of each other along this line, the proportion of $v_{d}$ will increase. Consequently, as the number of hypercubes increases, the complexity will significantly decrease.
In the case of a simple arrangement, in which every $d$ hyperplane in the arrangement form a vertex and every vertex is non-degenerate (see the RHS of Figure 8, $v=\binom{n}{d},v_{d}\geq\binom{n-d}{d}$), then Eq.(8) can be simplified to
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})$ $=\mathcal{O}\left(n^{2}d^{2}\left(\binom{n}{d}-\binom{n-d}{d}\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(\frac{n^{d}-(n-d)^{d}}{d!}+o(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(\frac{d^{2}n^{d-1}}{d!}+o(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(d\frac{n^{d-1}}{(d-1)!}+o(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
@@
In a specific arrangement, because the proportion of $v_{d}$ is different, the overall computational complexity needs to be analyzed separately. But the upper bound of the computational complexity is $\mathcal{O}(n^{2}d^{2}v)$, which is significantly smaller than $\mathcal{O}(nd^{2}v^{2})$ in Moss [25]. In brief, the algorithm developed by Avis and Fukuda [4] cannot guarantee enumerating all vertices, and the algorithm by Moss [25] costs $\mathcal{O}(nd^{2}v^{2})$, while our algorithm cost at most $\mathcal{O}(n^{2}d^{2}v)$. We reasonably contend that our algorithm is state-of-the-art in the problem of VE of an arrangement.
The lower bound of the computational complexity of the VE algorithm is $\mathcal{O}(ndv)$, since at least given each vertex, one needs to take $\mathcal{O}(nd)$ to obtain another dictionary. Our algorithm achieves this lower bound at $v_{d}$ vertices. This is because the valid reverse examination is no longer needed for vertices in the $d$-th layer. The Zero rule can guarantee the termination in $d$ steps. Can we further reduce the computational complexity of the proposed algorithm? It is difficult. To do so, one may seek a new pivot rule that either increases the portion of dictionaries that waive a valid reverse examination or reduces the number of examinations per valid reverse test. In the first situation, the potential superior rule needs to have pivot steps in $\mathcal{O}(\log(d))$ or $\mathcal{O}(1)$ to terminate. However, this requirement exceeds the capabilities of polynomial pivot rules, as there is always a certain dictionary that terminates in at least $d$ steps i.e., those dictionaries discussed in Proposition 5. In the second situation, reducing the examination means less information from a dictionary but more pivot steps. Given these challenges, the proposed Zero rule is a good choice for the vertex enumeration of an arrangement.
Figure 9: A representative example to show the superiority of the Zero rule over the Criss-Cross rule.
5 Example
In this section, we provide a simple yet representative example to demonstrate the superiority of using the Zero rule. This example exposes the weaknesses of the Criss-Cross rule: First, it contains a degenerate point that will make the VE algorithm with the Criss-Cross rule go into loops. Second, the Criss-Cross rule cannot guarantee the convergence in $d$ steps. Third, the Criss-Cross rule’s convergence is not unique, which forms multiple search trees. In contrast, the VE algorithm with the Zero rule is free of loops and can find all vertices. It also only forms a single tree with the depth of $d$. In this example, we slightly abuse the dictionary and the equation between slack variables interchangeably for convenience. It is worth noting that, even though the example we have chosen demonstrates many drawbacks of the original algorithm, it is not particularly special. The coefficients of the hyperplane and their corresponding constant terms are all randomly generated, and many of the properties reflected in this example can also be found in others.
Let us consider the arrangement in Figure 9, which has 5 hyperplanes, and the coefficients and biases of each hyperplane are
@@ 5 Example
and slack variables are $x_{1},...,x_{5}$:
$\begin{cases}&x_{1}=4-y_{1}-3y_{2}\\ &x_{2}=5-5y_{1}-y_{2}\\ &x_{3}=2-3y_{1}-2y_{2}\\ &x_{4}=1+y_{1}+3y_{2}\\ &x_{5}=-\frac{1}{2}+2y_{1}-y_{2}.\end{cases}$
The Zero rule: Taking $x_{1}$ and $x_{2}$ as co-basis and adding the variable $x_{g}$, we obtain the following initial dictionary:
@@
The dictionary with $N=\{g,1,2\}$ holds the lexicographically smallest order among all co-basis in this arrangement. However, this dictionary fails the lex-min test, which means we will not use this dictionary to produce a vertex.
Next, we examine each entry and select the one with the smallest index that satisfies the condition of the valid reverse. $(3,1)$ in Eq. (10) is not a valid reverse, while $(4,1)$ is. Therefore, we pivot $(4,1)$ and obtain the dictionary below:
@@
This dictionary also fails the lex-min test. We need to continue testing reverses based on this dictionary. $(1,2)$ is invalid, and $(3,2)$ is valid. Therefore, we pivot (3,2) and obtain the following dictionary:
@@
This dictionary successfully passes the lex-min test. Therefore, we can output the basis $B=\{1,2,5\}$. The current dictionary is at the $2$-nd layer, and there is no need to proceed further. As a result, we can skip checking the valid reverses from this dictionary and pivot (2,3) to return to Eq. (11).
After returning to Eq. (11), we only need to search entries whose row numbers or column numbers are larger than 3. While $(5,2)$, $(1,4)$, $(3,4)$, $(5,4)$ are all invalid reverses, Eq. (11) is finished. We can apply the Zero rule on it and go to the dictionary Eq. (10). Next, we continue testing and repeat this process. The LHS of Figure 10 presents a spanning tree that encompasses all dictionaries and their orders in the reverse process using the Zero rule.
Figure 10: Spanning trees that connect all vertices obtained by the VE algorithm. LHS: Zero rule. RHS: the Criss-Cross rule.
The Criss-Cross rule: Now, we utilize the Criss-Cross rule to solve this small example. Note that starting from any dictionary, repetitively applying the Criss-Cross rule will eventually lead to an optimal dictionary. Thus, the initial dictionary chosen for the VE algorithm is considered as an optimal dictionary. However, the optimal dictionary is not unique. In this example, we set the initial dictionary as Eq. (13), which is the same as the Zero rule.
@@
Meanwhile, another optimal dictionary is
@@
We need to apply the Zero rule based on these two dictionaries, respectively. Let us start from Eq. (13). Similarly, this dictionary does not pass the lex-min test, and $(4,1)$ is a valid reverse pivot about the Criss-Cross rule. Pivoting $(4,1)$ can obtain the following dictionary:
$\begin{cases}x_{1}&=5x_{g}+0x_{2}-x_{4}\\ x_{3}&=0x_{g}-\frac{1}{2}x_{2}+\frac{1}{2}x_{4}\\ x_{5}&=\frac{5}{2}x_{g}-x_{2}-\frac{1}{2}x_{4}\\ x_{f}&=-5x_{g}-0x_{2}+2x_{4}.\end{cases}$ (15)
which does not pass the lex-min test, either.
In Eq. (15), $(5,2)$ is a valid reverse pivot according to the Criss-Cross rule. Pivoting it results in the dictionary shown in Eq. (16). This new dictionary successfully passes the lex-min test, and we can compute the vertex from it.
$\begin{cases}x_{1}&=5x_{g}-x_{4}+0x_{5}\\ x_{2}&=5x_{g}-x_{4}-2x_{5}\\ x_{3}&=-\frac{5}{2}x_{g}+x_{4}+x_{5}\\ x_{f}&=-10x_{g}+2x_{4}+2x_{5}.\end{cases}$ (16)
Upon examining Eq. (16), we find that there are no valid reverses according to the Criss-Cross rule. Next, we apply the Criss-Cross rule to this dictionary, which identifies $(3,4)$. However, this step initiates an endless cycle in the VE algorithm, involving dictionaries with co-basis $\{4,5\}$, $\{3,5\}$, $\{1,3\}$, $\{1,2\}$, and $\{1,5\}$. The cycle is depicted in the RHS of Figure 10.
Next, let us apply the Criss-Cross rule on Eq. (14). No entries are valid reverses; therefore, Eq. (14) is an isolated dictionary.

### v2

@@ 4.3 Complexity Analysis
Given a hyperplane arrangement in $\mathbb{R}^{d}$ consisting of $n$ distinct hyperplanes with $v$ vertices, there are $\mathcal{O}(v)$ dictionaries. We note that applying the proposed Zero pivot takes at most $d$ iterations. Therefore, we denote the number of dictionaries that require exactly $d$ iterations as $\mathcal{O}(v_{d})$. Note that when calculating a new dictionary, one needs $(n-d)(d+1)=\mathcal{O}(nd)$ calculations. Testing the lex-min also needs $\mathcal{O}(nd)$ calculations.
4.3.1 Complexity of The AF and The Enhanced AF Algorithm
The computational complexity of the Enhanced AF Algorithm consist of the lex-min test on each dictionary, checking and performing the valid reverse pivot, and performing the Criss-Cross pivot. The total operation required for checking and performing the valid reverse pivot can be found in Appendix, Eq. (17), which is $\mathcal{O}(n^{2}dv)$. The lex-min tests and the Criss-Cross pivot together require at most $\mathcal{O}(ndv+(n+nd)v)=\mathcal{O}(ndv)$ operations. Therefore, its computational complexity is
$\mathcal{O}(ndv+n^{2}dv)=\mathcal{O}(n^{2}dv).$
As for the space complexity, since there is no additional storage, we only need to store those variables generated during the computation, which is in the order of $\mathcal{O}(nd)$.
Before proceeding with the analysis of the AF algorithm, it is important to note that the algorithm aims to maintain a balanced efficiency among all hyperplane arrangements. Therefore, we need to analyze the general case. Thus, we can assume that the probability of any position fulfilling the examination condition in each dictionary is $0.5$, regardless of the necessary condition in the original AF algorithm or the if-and-only-if condition used in the Enhanced AF algorithm. To facilitate the analysis in the following paragraph, we state the necessary condition in [2] as follows:
Proposition 6.
If $(s,r),s\in B_{\neq f},r\in N_{g}$, is a valid reverse criss-cross pivot for a dictionary $x_{B}=Ax_{N}$, then either:
$(a)$ $~{}\bar{a}_{sg}>0,\bar{a}_{sr}>0,\bar{a}_{sj}\geq 0\ for\ j\in N_{\neq g},j<s,$
$(b)$ $~{}\bar{a}_{fr}<0,\bar{a}_{sr}<0,\bar{a}_{ir}\leq 0\ for\ i\in B_{\neq f},i<r.$
Note that the AF algorithm requires processing each position given every dictionary exactly once. This means that checking and performing the valid reverse pivot in this algorithm requires $\mathcal{O}(ndv)$ times processing in the LHS of Figure 5, where there are $\mathcal{O}(v)$ valid reverse pivots.
We can categorize $\mathcal{O}(ndv)$ pivot check into three types: valid reverse pivots, those meet the necessary condition but are invalid, and those that fail the necessary condition. For the first two types, their cost is $\mathcal{O}(nd+n+n)=\mathcal{O}(nd)$. Observing Proposition 8 in Appendix and the necessary condition, they require checking at most $2(2d+2(n-d))=4n$ positions and $d+(n-d)=n$ positions, respectively. Thus, the second type occurs with a probability of $2^{-n+d-1}+2^{-d-1}-2^{-2n-1}-2^{-2n-1}=2^{-n+d-1}+2^{-d-1}-2^{-2n}$. Therefore, the total cost for the first two types combined is
$\mathcal{O}\Big{(}ndv+(2^{-n+d-1}+2^{-d-1}-2^{-2n})n^{2}d^{2}v\Big{)}$
$=$ $\mathcal{O}\Big{(}ndv+(2^{-n+d}+2^{-d})n^{2}d^{2}v\Big{)}.$
The third type incurs a cost of
$\mathcal{O}\Big{(}(ndv-v-(2^{-n+d-1}+2^{-d-1}-2^{-2n})ndv)n\Big{)}$
$=$ $\mathcal{O}\Big{(}(nd-1-(2^{-n+d}+2^{-d})nd)nv\Big{)}$
$=$ $\mathcal{O}\Big{(}(1-2^{-n+d}-2^{-d})n^{2}dv\Big{)}$
operations. Thus, the total operations in the AF algorithm requires for checking and performing valid reverse pivots are
$\mathcal{O}\Big{(}ndv+(2^{-n+d}+2^{-d})n^{2}d^{2}v+(1-2^{-n+d}-2^{-d})n^{2}dv\Big{)}$ (8)
$=$ $\mathcal{O}\Big{(}ndv+(2^{-n+d}+2^{-d})(d-1)n^{2}dv+n^{2}dv\Big{)}$
$=$ $\mathcal{O}\Big{(}(2^{-n+d}+2^{-d})n^{2}d^{2}v+n^{2}dv\Big{)}.$
In addition to the computations in Eq. (8), the consuming of those lex-min tests and those Criss-Cross pivots also require at most $\mathcal{O}(ndv+(n+nd)v)=\mathcal{O}(ndv)$ operations. Therefore, its total computational complexity should be
$\mathcal{O}\Big{(}(2^{-n+d}+2^{-d})n^{2}d^{2}v+n^{2}dv+ndv\Big{)}=\mathcal{O}(n^{2}dv).$
Similar with the analysis in the Enhanced AF algorithm, it needs to apply the reverse search to each optimal dictionary, and thus the space complexity should be $\mathcal{O}(nd)$.
Although this computational complexity is the same as that of Enhanced AF, we have omitted many lower-order terms compared to Enhanced AF. Moreover, the $(2^{-n+d}+2^{-d})n^{2}d^{2}v$ could be negligible only when the dimension is sufficiently large and the number of hyperplanes should be even larger, which is quite difficult to achieve in practice. Therefore, in our empirical experiments, the Enhanced AF performs better.
4.3.2 Complexity of The Moss Algorithm
@@
Now, let us analyze the computational complexity of the method from Moss [29]. In each dictionary, one needs to compute the ratio of the column of $x_{g}$ vs every column and then select the extreme value. This step incurs a cost of $\mathcal{O}(d(n-d)v)=\mathcal{O}(ndv)$. Subsequently, during the comparison with discovered dictionaries, their co-basis needs to be checked. If there are $k$ recorded dictionaries, then the comparison requires $\mathcal{O}(kd)$ for each candidate. When generating the adjacent co-basis about a fixed dictionary, one may find at most $nd$ candidates from it. Thus, comparing each candidate takes $\mathcal{O}(knd^{2})$ computations. Since there are a total of $\mathcal{O}(v)$ dictionaries, the total amount of computations is $\mathcal{O}(\sum_{k=1}^{v}(knd^{2}))=\mathcal{O}(nd^{2}v^{2})$. It is important to note that one only needs to compute those dictionaries whose co-basis are not duplicated, which costs $\mathcal{O}(ndv)$ in computing dictionaries and determining if they pass the lex-min test. Consequently, the complexity of the method from Moss [29] in VE is $\mathcal{O}(ndv+nd^{2}v^{2}+ndv)=\mathcal{O}(nd^{2}v^{2})$.
As for the space complexity, since the Moss algorithm needs to store and study each dictionary in the arrangement, the space complexity should be $\mathcal{O}(ndv)$.
4.3.3 Complexity of Our Algorithm
@@
It is important to note that each dictionary needs to be computed at most twice, namely through the valid reverse Zero pivot to enter it and through the Zero pivot to leave it. Therefore, the total cost of computing those dictionaries should be $\mathcal{O}(ndv)$. For each entry of a dictionary, testing a valid reverse Zero pivot requires at most $\frac{1}{2}(n-d+n)d$ examinations, which is of $\mathcal{O}(nd)$. While $\mathcal{O}(v-v_{d})$ dictionaries need to be tested for validity, $\mathcal{O}(v_{d})$ dictionaries do not. Thus, the total complexity incurred by testing validity is $\mathcal{O}(n^{2}d^{2}(v-v_{d}))$. In addition, each selection in the Zero rule also brings $\mathcal{O}(nd)$ comparisons, and we need to apply the selection on $\mathcal{O}(v-v_{d})$ dictionaries. Together with the operations conducted in the lex-min test, the complexity of our algorithm is approximately
$\mathcal{O}\Big{(}ndv+n^{2}d^{2}(v-v_{d})+ndv+ndv\Big{)}=\mathcal{O}\Big{(}n^{2}d^{2}(v-v_{d})+ndv_{d}\Big{)}$ (9)
When it comes to the space complexity, since there is no additional storage, we only need to store those variables generated during the computation, which can take the space of at most two dictionaries, in the order of $\mathcal{O}(nd)$. We organize the complexity of our algorithm into a theorem:
Theorem 1.
Given a hyperplane arrangement in $\mathbb{R}^{d}$ consisting of $n$ distinct hyperplanes and $v$ vertices, let $\mathcal{O}(v_{d})$ denote the number of dictionaries that require exactly $d$ steps of the Zero pivot. Then, the computational and space complexity of our algorithm is $\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})$ and $\mathcal{O}(nd)$, respectively.
@@
The ratio of $v_{d}$ vs $v$ has a significant impact on the actual complexity. Below we use two typical classes of hyperplane arrangements to illustrate the effect of $v_{d}$: one is a special case that consists of a lot of parallel hyperplanes; the other is a simple arrangement. Before estimation, we first characterize what kinds of dictionaries lie in the $d$-th layer.
@@
In an arrangement with a lot of parallel hyperplanes, where hyperplanes form boundaries of hypercubes arranged along a specific coordinate axis, such as $y_{1},y_{2},\ldots,y_{d}=0,1,y_{1}=2,y_{1}=3,y_{1}=4,\ldots,y_{1}=n-2d+1$. As shown in the LHS of Figure 7, $v_{d}=2^{-d}v$, then Eq. (9) can be simplified to
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})=\mathcal{O}((1-2^{-d})n^{2}d^{2}v+2^{-d}ndv)=\mathcal{O}(n^{2}d^{2}v).$
If one stacks multiple layers of hypercubes on top of each other along this line, the proportion of $v_{d}$ will increase. Consequently, as the number of hypercubes increases, the complexity for this class of hyperplane arrangements will significantly decrease.
In the case of a simple arrangement (see the RHS of Figure 7, $v=\binom{n}{d},v_{d}\geq\binom{n-d}{d}$), then Eq.(9) can be reduced to
$\mathcal{O}(n^{2}d^{2}(v-v_{d})+ndv_{d})$ $\leq\mathcal{O}\left(n^{2}d^{2}\left(\binom{n}{d}-\binom{n-d}{d}\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(\frac{n^{d}-(n-d)^{d}}{d!}+\mathcal{O}(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(\frac{d^{2}n^{d-1}}{d!}+\mathcal{O}(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
$=\mathcal{O}\left(n^{2}d^{2}\left(d\frac{n^{d-1}}{(d-1)!}+\mathcal{O}(n^{d-2})\right)+nd\binom{n-d}{d}\right)$
@@
In fact, the simple arrangement is a quite general case. If a hyperplane arrangement is randomly generated, it is very likely to be a simple arrangement, and therefore the complexity of our algorithm is significantly smaller than its counterparts. We reasonably contend that our algorithm is state-of-the-art in the problem of VE of an arrangement.
The lower bound of the computational complexity of the reversal VE algorithm is $\mathcal{O}(ndv)$, since at least given each vertex, one needs to take $\mathcal{O}(nd)$ to obtain another dictionary. Our algorithm achieves this lower bound at $v_{d}$ vertices. This is because the valid reverse examination is no longer needed for vertices in the $d$-th layer. The Zero rule can guarantee the termination in $d$ steps. Can we further reduce the computational complexity of the proposed algorithm? It is difficult. To do so, one may seek a new pivot rule that either increases the portion of dictionaries that waive a valid reverse examination or reduces the number of examinations per valid reverse test. In the first situation, the potential superior rule needs to have pivot steps in $\mathcal{O}(\log(d))$ or $\mathcal{O}(1)$ to terminate. However, this requirement is impossible, as there is always a certain dictionary that terminates in at least $d$ steps i.e., those dictionaries discussed in Proposition 5. In the second situation, reducing the examination means less information from a dictionary but more pivot steps. Given these challenges, the proposed Zero rule is a good choice in terms of achieving the lower bound over at least $v_{d}$ vertices.
5 Toy Example
Figure 8: Representative example.
In this section, we present a simple yet illustrative example to demonstrate the advantages of employing the Zero rule. This example highlights the limitations of employing the Criss-Cross rule: firstly, the Criss-Cross pivot does not guarantee convergence within $d$ steps. Secondly, regardless of the chosen objective function, there may exist multiple optimal dictionaries, resulting in multiple search trees. It is noteworthy that the example here is not particularly special, since the weights and biases of the hyperplane are generated randomly.
Let us consider the arrangement in Figure 8, which has 5 hyperplanes, and the weights and biases of each hyperplane are
@@ 5 Toy Example
and slack variables are $x_{1},...,x_{5}$, which are expressed in the following equations:
$\begin{cases}&x_{1}=4x_{g}-y_{1}-3y_{2}\\ &x_{2}=5x_{g}-5y_{1}-y_{2}\\ &x_{3}=2x_{g}-3y_{1}-2y_{2}\\ &x_{4}=x_{g}+y_{1}+3y_{2}\\ &x_{5}=-\frac{1}{2}x_{g}+2y_{1}-y_{2}\end{cases}$
Ours: Taking $x_{1}$ and $x_{2}$ as co-basis and adding the variable $x_{g}$, we obtain the following initial dictionary:
@@
The dictionary with $N=\{g,1,2\}$ is the lexicographically minimal among all co-basis in this arrangement. We assign it with $\mathtt{layer}=0$. As for the lex-min test, for convenience, denote this dictionary as $(B^{0},N^{0},\bar{A}^{0})$. Since there exists a pair of variables $(5,1)$, where $5\in B^{0}$ and $1\in N^{0}$, such that $\bar{a}_{3g}^{0}=0$ and $\bar{a}_{32}^{0}=\frac{1}{2}\neq 0$, $B^{0}$ is not lexicographically minimal for the corresponding basic solution. Therefore, it did not pass the lex-min test, and we do not need to print its basis.
Next, we examine each entry sequentially, pivot $(3,1)$ on the dictionary (11) is not a valid reverse Zero pivot, while $(4,1)$ is. Note that the layer of dictionary (11) is 00, which is less than $d-1=1$, then we pivot $(4,1)$ and obtain a dictionary in $\mathtt{layer}=0+1=1$, and update the dictionary as below:
@@
Now we do the lex-min test for this dictionary. Denote this dictionary as $(B^{1},N^{1},\bar{A}^{1})$, we see there exists a pair of variables $(3,2)$, where $3\in B^{1}$ and $2\in N^{1}$, such that $\bar{a}_{3g}^{1}=0$ and $\bar{a}_{32}^{1}=\frac{1}{2}\neq 0$. According to Proposition 5, $B^{1}$ is not lexicographically minimum for the corresponding basic solution. Therefore, it does not pass the lex-min test, and we do not print its basis.
After the lex-min test, we resume the reverse search in this dictionary. The pivot $(1,2)$ is invalid, and the pivot $(3,2)$ is valid. Note that the layer of dictionary (12) is $1$, which is equal to $d-1=1$, then we pivot $(3,2)$ and obtain the following dictionary in $\mathtt{layer}=1+1=2$ temporarily but do not update the dictionary being examined:
@@
Denote the dictionary by $(B^{temp},N^{temp},\bar{A}^{temp})$, the only variable $r\in B^{temp}$ such that $\bar{a}^{temp}_{rg}=0$ is $2$. Since $\forall s\in N_{\neq g},2<s$, Proposition 5 implies this $B^{temp}$ attain the lexicographically minimum for the corresponding basic solution; therefore, we output its basis $\{1,2,5\}$.
Consequently, we continue searching on the dictionary being examined, which is the dictionary (12). After checking all entries on this dictionary, all remaining entries are invalid reverses, the dictionary (12) is finished. Then, we can apply the Zero rule on it to go to the dictionary (11) and so forth. The LHS of Figure 9 presents a spanning tree that encompasses all dictionaries and their orders in our algorithm.
Figure 9: Spanning trees that connect all vertices obtained by the VE algorithm. LHS: Ours. RHS: Enhanced AF.
Enhanced AF: Now, we utilize the Enhanced AF algorithm to solve this small example. The first step is to construct an optimal dictionary with a proper objective function. According to [2], we set the initial dictionary as the dictionary (14), which is adding a row of objective function to dictionary (11).
@@
Meanwhile, another optimal dictionary in the hyperplane arrangement is
@@
The algorithm needs to to apply the reverse search about the Criss-Cross rule on them, respectively. Let us start from the dictionary (14).
Similar with the analysis to the dictionary (11), this dictionary does not pass the lex-min test, and the pivot $(3,1)$ is a valid reverse Criss-Cross pivot. Pivoting $(3,1)$ can obtain the following dictionary:
$\begin{cases}x_{1}&=5x_{g}-x_{2}-2x_{3}\\ x_{4}&=0x_{g}-x_{2}+2x_{3}\\ x_{5}&=\frac{5}{2}x_{g}-x_{2}-x_{3}\\ x_{f}&=-5x_{g}+0x_{2}+2x_{3}.\end{cases}$ (16)
Denote this dictionary by $(B^{2},N^{2},\bar{A}^{2})$, then there exists a pair of variables $(4,2),4\in B_{\neq f}^{2},2\in N^{2}$ such that $\bar{a}^{2}_{4g}=0$ and $\bar{a}^{2}_{42}=-1\neq 0$. It implies that the dictionary fails the lex-min test.
After carefully checking all entries of this dictionary, all remaining entries are invalid reverses, and the study on dictionary (16) is finished. We can apply the Criss-Cross rule on it and go to the dictionary (14). Next, we continue the test for the pair of variables after $(3,1)$ and so forth.
On the other hand, we apply the reverse search with the Criss-Cross rule on the dictionary (15). There is no valid reverse Criss-Cross pivot on it; therefore, the dictionary (15) is an isolated dictionary. Thus, the entire VE algorithm terminated. The RHS of Figure 9 presents a spanning tree that encompasses all dictionaries and their orders by Enhanced AF.